ericas age is 8 and her mother is 42. in how many years time will the mother be 3 times older than her daughter?
Answer by fcabanski(1391)

Add some unknown number x years to both ages. Erica will be 8+x years old. Mom will be 42 + x. The goal is to find x so that 3(8+x) = 42 + x - mom is three times as old.


3(8+x) = 42 + x


24 + 3x = 42 + x


2z = 18


x=9


In 9 years Erica will be 17 and mom will be 51. 3 x 17 = 51